OpenStudy (anonymous):

Is the relationship shown by the data linear? If so, model the data with an equation. x y –9 –2 –5 –7 –1 –12 3 –17 The relationship is linear; y + 2 = 4/5(x + 9). The relationship is linear; y + 2 = -4/5(x + 9). The relationship is linear; y + 9 = -5/4(x + 2). The relationship is not linear.

OpenStudy (anonymous):

@LynFran

OpenStudy (anonymous):

sebaxtiangb

OpenStudy (anonymous):

@sebaxtiangb

OpenStudy (anonymous):

not sure what to do

OpenStudy (sebaxtiangb):

i think is not linear

OpenStudy (anonymous):

ok. Can you help me with other ones?

OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. 2x + 9y = –30; (–8, 0) y = 2/9x + 36 y = 9/2x + 36 y = -9/2x + 36 y = 2/9x – 8

OpenStudy (sebaxtiangb):

this is flvs?

OpenStudy (anonymous):

What is flvs?

OpenStudy (anonymous):

This is integrated math

OpenStudy (anonymous):

@sebaxtiangb you there?

OpenStudy (sebaxtiangb):

sorry

OpenStudy (sebaxtiangb):

i do flvs sorry

OpenStudy (anonymous):

What is flvs?

OpenStudy (sebaxtiangb):

florida virtual schools

OpenStudy (anonymous):

ok. Can you still help me?

OpenStudy (sebaxtiangb):

i dont know that

OpenStudy (anonymous):

ok. same here. I try another one.

OpenStudy (anonymous):

Write an equation for each translation of y=|x|. 5.5 units right y = | x | – 5.5 y = | x – 5.5 | y = | x + 5.5 | y = | x | + 5.5

OpenStudy (sebaxtiangb):

is the absolute value of the hole thing

OpenStudy (anonymous):

it a different question.Don't now how to put it into an equation